A projectile starting at a velocity of 25m/s from ground

hits a target on the ground located at a distance of 1000

meters after 40 seconds.

a) What is the size of the angle ??

b) What is the magnitude of the velocity of the object

just before it hits the target?



1
Expert's answer
2021-08-12T06:19:34-0400


Find the horizontal component using the range and time:


"v_x=\frac Rt=\frac{1000}{40}=25\text{ m/s}."


The vertical velocity as the body hits the target is


"v_y=g\frac{t}{2}=196\text{ m/s}."

The angle:


"\alpha=\arctan\frac{196}{25}=82.7°."

The velocity just before the body hits the target:


"v=\sqrt{v_x+v_y}=197.6\text{ m/s}."